OpenStudy (anonymous):

18x^2-5x=0 I'm having trouble factoring. That or i completely don't know how to do this.

OpenStudy (jonnyvonny):

From this equation, what can you factor out?

OpenStudy (anonymous):

the x and get 18x-5

OpenStudy (jonnyvonny):

Yes, but don't neglect the x; it has feelings too. x(18x-5)=0. Is x=0 the only solution?

OpenStudy (jonnyvonny):

Remember, a solution is when, after putting an x-value, the function would equal 0. We know that when x=0, then, since we're multiplying, regardless of what the value of 18x-5 would be, the outcome would be 0.

OpenStudy (jonnyvonny):

0*(18(0)-5)=0, since this is true, can we not find when 18x-5 would equal 0 aswell?

OpenStudy (anonymous):

is 3x(x+2)=9 an easier example? I seem to do better when I see it laid out step by step instead of words. I get confused very easy

OpenStudy (jonnyvonny):

Icy icy, people learn differently, also, it is math, so seeing it IS easier. so..

OpenStudy (jonnyvonny):

The 9 we should move to the other side, so we may find when the function would equal 0.\[3x(2x+2)=9\rightarrow 3x(2x+2)-9=0\]

OpenStudy (anonymous):

ok but why is it 2x+2?

OpenStudy (jonnyvonny):

It is given; its what you posted. Everytime you get a new math problem, everything that was revalent to it is tossed out, the intercepts, the slope, the vertex, everything. The reason why we want 2x+2=0 is because when 2x+2=0, the entire rest will equal 0.

OpenStudy (anonymous):

oh i get it now!

OpenStudy (jonnyvonny):

That's great! Excellent! What made it click?